The density of mercury is 13.6 g/cm. Estimate the b value

Grade:11

2 Answers

Tanvi
6 years ago
Consider 1 mole of mercury - that`s about 200g. Find its volume. This is the volume occupied by 1 mole of mercury atoms. So that is (roughly) the value of b.
Gauri
3 years ago
b= 4 * volume (for one molecule)
b= 4 * molecular mass/ density
= 4*200/13.6 = 800/13.6 = 58.82 cm^3/mole